# Temperature Scaling Attack Disrupting Model Confidence in Federated Learning

### v1

@@ 1 Introduction
Predictive confidence is a cornerstone of operational reliability in modern safety-critical systems [park2023self]. In domains such as autonomous driving [antonante2023task, kendall2017uncertainties], remote clinical monitoring [park2020heartquake], and industrial robotics [lee2026iros], probability outputs are not merely explanatory; they are consumed as foundational control signals for risk-aware policies [kompa2021second]. For example, a vehicle may engage conservative fallback behaviors when confidence is low, and a healthcare pipeline may escalate uncertain cases for human review [geifman2017selective]. These mechanisms implicitly assume calibration: a confidence score should correspond to the empirical likelihood of correctness [niculescu2005predicting]. When this alignment is systematically distorted, over-confidence can suppress vital safeguards, while under-confidence can trigger excessive verification, degrading both safety and operational efficiency.
@@
In many such systems, federated learning (FL) is increasingly adopted to reduce the need for centralizing sensitive user data while enabling personalization at scale. However, decentralization fundamentally enlarges the adversarial surface. Existing FL security research has largely remained accuracy-centric, focusing on poisoning attacks that disrupt convergence or backdoors that induce targeted misclassifications [fang2020local, bagdasaryan2020backdoor, xie2019dba]. While effective at causing errors, these attacks can often be detectable in practice [lee2024tazza, blanchard2017machine, lee2024detrigger]; noticeable drops in validation accuracy or abnormal training dynamics serve as immediate red flags, triggering server-side auditing and mitigation.
@@
For a sophisticated adversary, a more insidious threat is to preserve model utility while subverting its operational reliability [zhu2022rethinking]. By targeting to attack the calibration rather than accuracy, the attacker can keep the model deployable and continue to deliver seemingly correct predictions to end users, while systematically distorting the probability scale that downstream policies rely on [obadinma2024calibration]. This creates a silent failure mode, given that deployed systems typically monitor only accuracy or loss, as calibration assessment requires representative validation data that is rarely available post-deployment. Consequently, the system may pass standard accuracy-based validation while safety gating and triage logic driven by model confidence become systematically unreliable.
@@
Such a risk is amplified in federated settings with inherently non-IID client data [mcmahan2017communication]. Post-hoc calibration is also challenging to apply: each client’s validation set is small, biased, and often mismatched to the global deployment distribution [chen2024watch, peng2024fedcal]. As a result, local corrections are weakened and “normal” confidence behavior becomes ill-defined across clients and rounds, hindering reliable auditing of calibration manipulation.
@@
2 Background and Related Work
2.1 Security Threats in Federated Learning
@@
(i) Untargeted model poisoning. Adversaries perturb local updates to degrade global accuracy or slow convergence, using techniques such as sign flipping, scaled gradients, or noise injection [bhagoji2019analyzing, lee2024tazza]. These attacks aim to disrupt the learning process and render the global model unusable. (ii) Backdoor and targeted poisoning. Attackers train on poisoned data so that inputs containing specific triggers are mapped to attacker-chosen labels while maintaining accuracy on clean data [bagdasaryan2020backdoor, xie2019dba, lee2024detrigger]. Model replacement attacks amplify this by scaling malicious updates to dominate the aggregation, effectively overwriting the global model with a backdoored version.
Byzantine-robust defenses and evasion. To counter malicious updates, robust aggregation mechanisms have been proposed, including Trimmed Mean [yin2018byzantine], Krum [blanchard2017machine], FLAME [nguyen2022flame], and FLTrust [cao2020fltrust]. In response, sophisticated adversaries have developed collusion strategies to evade these defenses by carefully shaping their updates to appear benign during the aggregation process [fang2020local].
@@
2.2 Model Confidence and Calibration
@@
$p_{\theta}(y=c\mid x)=\frac{\exp(z_{c}(x))}{\sum_{k=1}^{C}\exp(z_{k}(x))},\quad c\in\{1,\ldots,C\}.$ (1)
@@
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|,$ (2)
where $\mathrm{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}a_{i}$ and $\mathrm{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$ denote per-bin empirical accuracy and average confidence, respectively [guo2017calibration]. While ECE measures the magnitude of miscalibration, it discards directionality. This distinction matters: overconfidence suppresses escalation despite incorrect predictions, while underconfidence triggers unnecessary verification. Signed-ECE (sECE) retains this direction [kim2024curved]:
$\mathrm{sECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\left(\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right),$ (3)
@@
$\mathrm{NLL}=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(y=y_{i}\mid x_{i}),$ (4)
@@
$\mathrm{Brier}=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\left(p_{\theta}(y=c\mid x_{i})-e(y_{i})_{c}\right)^{2}.$ (5)
@@
2.3 Attacking Confidence Calibration
Prior work on attacking model calibration has largely focused on inference-time adversarial examples that craft input perturbations to induce miscalibration or unreliable uncertainty estimates [obadinma2024calibration, galil2021disrupting, ledda2023adversarial]. These attacks typically modify inputs to alter the posterior distribution (e.g., confidence or entropy) without changing the predicted label, often requiring white-box access to model internals (architecture, parameters, logits) or repeated black-box queries [goodfellow2014explaining, madry2017towards, carlini2017towards]. While effective in controlled settings, such approaches face two fundamental challenges in real-world deployments.
First, they assume the attacker can compute or estimate sensitive model signals, which may be unavailable due to API restrictions, output rounding, rate limits, or privacy constraints. Second, even when perturbations can be generated, their effectiveness degrades substantially in physical deployments due to sensing noise, compression artifacts, domain shifts, and non-differentiable pre-processing [kurakin2018adversarial, eykholt2018robust, lee2025mind, guo2024invisible, lee2023exploiting]. These challenges are amplified in federated learning, where adversaries cannot control other clients’ inputs or consistently access the global model’s gradients at inference across devices. Consequently, inference-time calibration attacks do not naturally extend to FL threat models that operate through local training.
In contrast, training-time manipulation of model confidence has received limited attention. Most FL attacks focus on degrading accuracy, targeted misclassification, or implanting backdoors [bagdasaryan2020backdoor, fang2020local, lyu2022privacy], objectives orthogonal to calibration integrity that can often be detected via accuracy monitoring or misprediction patterns. Our work addresses this gap by treating calibration as a first-class attack objective in FL. We propose a training-time mechanism that systematically biases the model’s confidence through temperature-controlled local optimization, without requiring inference-time perturbations or access to other clients’ data.
2.4 Temperature Scaling
@@
While temperature has been most commonly used as an inference-time control knob in cases such as knowledge distillation [hinton2015distilling], post-hoc calibration [guo2017calibration] or probabilistic sampling [jang2016categorical, zhu2024hot, zhang2024edt], recent work has begun to study temperature as a training-time factor that can meaningfully alter learning dynamics. In particular, Agarwala et. al analyze how softmax temperature interacts with optimization and generalization under softmax-cross-entropy training, highlighting that temperature can effectively modulate the training behavior beyond merely reshaping probabilities at test time [agarwala2020temperature].
In parallel, emerging work has explored temperature scaling during federated local training to improve convergence and generalization [lee2026improving]. Our work extends this theoretical foundation from convex, IID settings to realistic non-convex, non-IID federated environments. Crucially, we identify the learning rate–temperature coupling as fundamental: maintaining a constant effective step size $\beta{=}\eta/\tau$ preserves benign-style convergence bounds under client drift while enabling systematic confidence distortion. This finding serves an adversarial purpose (i.e., evading detection while degrading calibration) rather than improving performance.
3 Threat Model
@@
Assumptions. We assume that attackers can control local training but cannot interfere with server-side aggregation or benign clients, and that the fraction of malicious participants is less than $50\%$ [fang2020local]. The adversary can not manipulate other clients’ training process or data, and has no control over any computation performed at the server.
4 Temperature Scaling Attack
@@
4.1 Rationale
Figure 1: Effect of training-time temperature scaling on logit updates and test-time confidence. Solid curves show the $\tau$-scaled sigmoid used to compute updates (: $\tau{<}1$, : $\tau{>}1$), while evaluation uses the standard $\tau{=}1$ mapping (- -), so updated logits ($\bullet$ $\rightarrow$ $\bullet$) can yield shifted probabilities and degraded calibration; note the different x-axis ranges.
@@ 4.2 Overall Workflow
4.3 Theoretical Convergence Analysis
@@
$\|\nabla f_{k}(u)-\nabla f_{k}(v)\|\leq L\|u-v\|,\qquad\forall u,v.$ (7)
@@
$\mathbb{E}_{\xi}[\mathbf{g}_{k}(\mathbf{\theta};\xi)]=\nabla f_{k}(\mathbf{\theta})\quad and$ (8)
$\mathbb{E}_{\xi}\big[\|\mathbf{g}_{k}(\mathbf{\theta};\xi)-\nabla f_{k}(\mathbf{\theta})\|^{2}\big]\leq\sigma^{2},\qquad\forall\mathbf{\theta}.$ (9)
@@
$\sum_{k=1}^{K}p_{k}\|\nabla f_{k}(\mathbf{\theta})-\nabla F(\mathbf{\theta})\|^{2}\leq\Gamma^{2}.$ (10)
@@
$\eta_{k}/\tau_{k}=\beta,\qquad\forall k\in[K].$ (11)
@@
$\|\nabla_{\theta}^{2}\ell_{\tau}(\theta;x,y)\|\leq\frac{C(x)}{\tau^{2}},$ (12)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big[\|\nabla F(\theta_{t})\|^{2}\big]\leq$ $\frac{2\big(F(\theta_{0})-F^{\star}\big)}{\beta ET}\;+\;\;\underbrace{\frac{\beta L\sigma^{2}}{M}}_{\text{variance}}$ (13)
@@ 5.1 Setup
Federated Learning Configuration. Unless otherwise specified, we adopt a standard synchronous FL setup using FedAvg. In each FL round, the server samples $5$ clients (10%) uniformly at random from a total of $50$ clients and broadcasts the current global model. Each selected client performs $5$ steps of local SGD on its private data and returns either model parameters or model updates to the server, which aggregates them using data-size weights $p_{k}$. To reflect realistic statistical heterogeneity, we construct non-IID client partitions using a Dirichlet split [hsu2019measuring] with concentration parameter $\alpha$=1.0 (smaller $\alpha$ indicates higher heterogeneity), otherwise specified. We report results averaged over $5$ independently selected random seeds and evaluate both predictive performance and calibration using accuracy and calibration metrics.
5.2 Overall Attack Performance
@@
Table 1: Global model accuracy and confidence metrics with varying temperature on MNIST-MLP configuration.
@@
Table 2: Global model accuracy and confidence metrics with varying temperature on CIFAR10-CNN configuration.
@@
Table 3: Global model accuracy and confidence metrics with varying temperature on CIFAR100-ResNet18 configuration.
@@
Figure 2: Confidence distribution and reliability diagrams under different training temperatures $\tau$ on CIFAR10–CNN.
@@
5.3 Benign-Aligned Updates
Figure 3: Global accuracy (left) and sECE (right) over rounds on CIFAR10–CNN for different training temperatures $\tau$. Accuracy remains similar across $\tau$, while sECE shifts systematically, indicating controllable under-/over-confidence.

### v2

@@ 1. Introduction
Predictive confidence is a cornerstone of operational reliability in modern safety-critical systems (Park et al., 2023). In domains such as autonomous driving (Antonante et al., 2023; Kendall and Gal, 2017), remote clinical monitoring (Park et al., 2020), and industrial robotics (Lee et al., 2026a), probability outputs are not merely explanatory; they are consumed as foundational control signals for risk-aware policies (Kompa et al., 2021). For example, a vehicle may engage conservative fallback behaviors when confidence is low, and a healthcare pipeline may escalate uncertain cases for human review (Geifman and El-Yaniv, 2017). These mechanisms implicitly assume calibration: a confidence score should correspond to the empirical likelihood of correctness (Niculescu-Mizil and Caruana, 2005). When this alignment is systematically distorted, over-confidence can suppress vital safeguards, while under-confidence can trigger excessive verification, degrading both safety and operational efficiency.
@@
In many such systems, federated learning (FL) is increasingly adopted to reduce the need for centralizing sensitive user data while enabling personalization at scale. However, decentralization fundamentally enlarges the adversarial surface. Existing FL security research has largely remained accuracy-centric, focusing on poisoning attacks that disrupt convergence or backdoors that induce targeted misclassifications (Fang et al., 2020; Bagdasaryan et al., 2020; Xie et al., 2019). While effective at causing errors, these attacks can often be detectable in practice (Lee et al., 2024a; Blanchard et al., 2017; Lee et al., 2024b); noticeable drops in validation accuracy or abnormal training dynamics serve as immediate red flags, triggering server-side auditing and mitigation.
@@
For a sophisticated adversary, a more insidious threat is to preserve model utility while subverting its operational reliability (Zhu et al., 2022). By targeting to attack the calibration rather than accuracy, the attacker can keep the model deployable and continue to deliver seemingly correct predictions to end users, while systematically distorting the probability scale that downstream policies rely on (Obadinma et al., 2024). This creates a silent failure mode, given that deployed systems typically monitor only accuracy or loss, as calibration assessment requires representative validation data that is rarely available post-deployment. Consequently, the system may pass standard accuracy-based validation while safety gating and triage logic driven by model confidence become systematically unreliable.
@@
Such a risk is amplified in federated settings with inherently non-IID client data (McMahan et al., 2017). Post-hoc calibration is also challenging to apply: each client’s validation set is small, biased, and often mismatched to the global deployment distribution (Chen et al., 2024; Peng et al., 2024). As a result, local corrections are weakened and “normal” confidence behavior becomes ill-defined across clients and rounds, hindering reliable auditing of calibration manipulation.
@@
2. Background and Related Work
2.1. Security Threats in Federated Learning
@@
(i) Untargeted model poisoning. Adversaries perturb local updates to degrade global accuracy or slow convergence, using techniques such as sign flipping, scaled gradients, or noise injection (Bhagoji et al., 2019; Lee et al., 2024a). These attacks aim to disrupt the learning process and render the global model unusable. (ii) Backdoor and targeted poisoning. Attackers train on poisoned data so that inputs containing specific triggers are mapped to attacker-chosen labels while maintaining accuracy on clean data (Bagdasaryan et al., 2020; Xie et al., 2019; Lee et al., 2024b). Model replacement attacks amplify this by scaling malicious updates to dominate the aggregation, effectively overwriting the global model with a backdoored version.
Byzantine-robust defenses and evasion. To counter malicious updates, robust aggregation mechanisms have been proposed, including Trimmed Mean (Yin et al., 2018), Krum (Blanchard et al., 2017), FLAME (Nguyen et al., 2022), and FLTrust (Cao et al., 2020). In response, sophisticated adversaries have developed collusion strategies to evade these defenses by carefully shaping their updates to appear benign during the aggregation process (Fang et al., 2020).
@@
2.2. Model Confidence and Calibration
@@
(1) $p_{\theta}(y=c\mid x)=\frac{\exp(z_{c}(x))}{\sum_{k=1}^{C}\exp(z_{k}(x))},\quad c\in\{1,\ldots,C\}.$
@@
(2) $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|,$
where $\mathrm{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}a_{i}$ and $\mathrm{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$ denote per-bin empirical accuracy and average confidence, respectively (Guo et al., 2017). While ECE measures the magnitude of miscalibration, it discards directionality. This distinction matters: overconfidence suppresses escalation despite incorrect predictions, while underconfidence triggers unnecessary verification. Signed-ECE (sECE) retains this direction (Kim et al., 2024):
(3) $\mathrm{sECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\left(\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right),$
@@
(4) $\mathrm{NLL}=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(y=y_{i}\mid x_{i}),$
@@
(5) $\mathrm{Brier}=\frac{1}{n}\sum_{i=1}^{n}\sum_{c=1}^{C}\left(p_{\theta}(y=c\mid x_{i})-e(y_{i})_{c}\right)^{2}.$
@@
2.3. Attacking Confidence Calibration
Prior work on attacking model calibration has largely focused on inference-time adversarial examples that craft input perturbations to induce miscalibration or unreliable uncertainty estimates (Obadinma et al., 2024; Galil and El-Yaniv, 2021; Ledda et al., 2023). These attacks typically modify inputs to alter the posterior distribution (e.g., confidence or entropy) without changing the predicted label, often requiring white-box access to model internals (architecture, parameters, logits) or repeated black-box queries (Goodfellow et al., 2014; Madry et al., 2017; Carlini and Wagner, 2017). While effective in controlled settings, such approaches face two fundamental challenges in real-world deployments.
First, they assume the attacker can compute or estimate sensitive model signals, which may be unavailable due to API restrictions, output rounding, rate limits, or privacy constraints. Second, even when perturbations can be generated, their effectiveness degrades substantially in physical deployments due to sensing noise, compression artifacts, domain shifts, and non-differentiable pre-processing (Kurakin et al., 2018; Eykholt et al., 2018; Lee et al., 2025; Guo et al., 2024; Lee et al., 2023). These challenges are amplified in federated learning, where adversaries cannot control other clients’ inputs or consistently access the global model’s gradients at inference across devices. Consequently, inference-time calibration attacks do not naturally extend to FL threat models that operate through local training.
In contrast, training-time manipulation of model confidence has received limited attention. Most FL attacks focus on degrading accuracy, targeted misclassification, or implanting backdoors (Bagdasaryan et al., 2020; Fang et al., 2020; Lyu et al., 2022), objectives orthogonal to calibration integrity that can often be detected via accuracy monitoring or misprediction patterns. Our work addresses this gap by treating calibration as a first-class attack objective in FL. We propose a training-time mechanism that systematically biases the model’s confidence through temperature-controlled local optimization, without requiring inference-time perturbations or access to other clients’ data.
2.4. Temperature Scaling
@@
While temperature has been most commonly used as an inference-time control knob in cases such as knowledge distillation (Hinton et al., 2015), post-hoc calibration (Guo et al., 2017) or probabilistic sampling (Jang et al., 2016; Zhu et al., 2024; Zhang et al., 2024), recent work has begun to study temperature as a training-time factor that can meaningfully alter learning dynamics. In particular, Agarwala et. al analyze how softmax temperature interacts with optimization and generalization under softmax-cross-entropy training, highlighting that temperature can effectively modulate the training behavior beyond merely reshaping probabilities at test time (Agarwala et al., 2020).
In parallel, emerging work has explored temperature scaling during federated local training to improve convergence and generalization (Lee et al., 2026b). Our work extends this theoretical foundation from convex, IID settings to realistic non-convex, non-IID federated environments. Crucially, we identify the learning rate–temperature coupling as fundamental: maintaining a constant effective step size $\beta{=}\eta/\tau$ preserves benign-style convergence bounds under client drift while enabling systematic confidence distortion. This finding serves an adversarial purpose (i.e., evading detection while degrading calibration) rather than improving performance.
3. Threat Model
@@
Assumptions. We assume that attackers can control local training but cannot interfere with server-side aggregation or benign clients, and that the fraction of malicious participants is less than $50\%$ (Fang et al., 2020). The adversary can not manipulate other clients’ training process or data, and has no control over any computation performed at the server.
4. Temperature Scaling Attack
@@
4.1. Rationale
Figure 1. Effect of training-time temperature scaling on logit updates and test-time confidence. Solid curves show the $\tau$-scaled sigmoid used to compute updates (: $\tau{<}1$, : $\tau{>}1$), while evaluation uses the standard $\tau{=}1$ mapping (- -), so updated logits ($\bullet$ $\rightarrow$ $\bullet$) can yield shifted probabilities and degraded calibration; note the different x-axis ranges.
@@ 4.2. Overall Workflow
4.3. Theoretical Convergence Analysis
@@
(7) $\|\nabla f_{k}(u)-\nabla f_{k}(v)\|\leq L\|u-v\|,\qquad\forall u,v.$
@@
(8) $\mathbb{E}_{\xi}[\mathbf{g}_{k}(\mathbf{\theta};\xi)]=\nabla f_{k}(\mathbf{\theta})\quad and$
(9) $\mathbb{E}_{\xi}\big[\|\mathbf{g}_{k}(\mathbf{\theta};\xi)-\nabla f_{k}(\mathbf{\theta})\|^{2}\big]\leq\sigma^{2},\qquad\forall\mathbf{\theta}.$
@@
(10) $\sum_{k=1}^{K}p_{k}\|\nabla f_{k}(\mathbf{\theta})-\nabla F(\mathbf{\theta})\|^{2}\leq\Gamma^{2}.$
@@
(11) $\eta_{k}/\tau_{k}=\beta,\qquad\forall k\in[K].$
@@
(12) $\|\nabla_{\theta}^{2}\ell_{\tau}(\theta;x,y)\|\leq\frac{C(x)}{\tau^{2}},$
@@
(13) $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big[\|\nabla F(\theta_{t})\|^{2}\big]\leq$ $\frac{2\big(F(\theta_{0})-F^{\star}\big)}{\beta ET}\;+\;\;\underbrace{\frac{\beta L\sigma^{2}}{M}}_{\text{variance}}$
@@ 5.1. Setup
Federated Learning Configuration. Unless otherwise specified, we adopt a standard synchronous FL setup using FedAvg. In each FL round, the server samples $5$ clients (10%) uniformly at random from a total of $50$ clients and broadcasts the current global model. Each selected client performs $5$ steps of local SGD on its private data and returns either model parameters or model updates to the server, which aggregates them using data-size weights $p_{k}$. To reflect realistic statistical heterogeneity, we construct non-IID client partitions using a Dirichlet split (Hsu et al., 2019) with concentration parameter $\alpha$=1.0 (smaller $\alpha$ indicates higher heterogeneity), otherwise specified. We report results averaged over $5$ independently selected random seeds and evaluate both predictive performance and calibration using accuracy and calibration metrics.
5.2. Overall Attack Performance
@@
Table 1. Global model accuracy and confidence metrics with varying temperature on MNIST-MLP configuration.
@@
Table 2. Global model accuracy and confidence metrics with varying temperature on CIFAR10-CNN configuration.
@@
Table 3. Global model accuracy and confidence metrics with varying temperature on CIFAR100-ResNet18 configuration.
@@
Figure 2. Confidence distribution and reliability diagrams under different training temperatures $\tau$ on CIFAR10–CNN.
@@
5.3. Benign-Aligned Updates
Figure 3. Global accuracy (left) and sECE (right) over rounds on CIFAR10–CNN for different training temperatures $\tau$. Accuracy remains similar across $\tau$, while sECE shifts systematically, indicating controllable under-/over-confidence.